# The Binary Fraction of Stars in the Dwarf Galaxy Ursa Minor via Dark Energy Spectroscopic Instrument

### v1

@@ 3 Methodology
Motivated by the method developed in 2017AJ....153..254S and 2018AJ....156..257S, we introduce several improvements to achieve a more realistic constraint on the binary fraction and a more robust estimation of its uncertainty. Specifically, we refine both the modeling of orbital parameters and the construction of the likelihood function. The detailed methodological differences will be described in the following subsections.
@@
3.1 Binary Orbital Parameters
@@
We adopt Equation 1 to simulate the LOSV at different epochs of a given star (see more details in Section 3.2 below). To achieve this goal, we have to know the probability distribution of the seven orbital parameters, i.e., $m_{1}$, $q$, $P$, $e$, $i$, $\omega$, and $\theta$.
Among these parameters, $q$, $m_{1}$, $e$, and $P$ are intrinsic, while the other three angles of $i$, $\theta$, and $\omega$ depend on the observer’s perspective. We adopt the distributions of the intrinsic parameters ($q$, $e$, $P$) from two different models developed by 1991A&A...248..485D(hereafter 1991A&A...248..485D) and 2017ApJS..230...15M(hereafter 2017ApJS..230...15M), which constrain the binary orbital parameter distributions based on solar neighborhood observations. The 1991A&A...248..485D model is widely used in many previous studies. It provides independent distributions for each orbital parameter and is based on solar-like stars, making it straightforward to apply. It is important to note that our member stars in UMi are all giant stars with metallicity of $\mathrm{[Fe/H]}<-1$, significantly lower than those in the solar neighborhood. This means that by adopting the 1991A&A...248..485D model, we are assuming that the binary properties of giants in dwarf galaxies are similar to those of solar-like stars in the solar neighborhood.
@@
We refer the readers to the original 1991A&A...248..485D and 2017ApJS..230...15M papers for more details, and in the following, we briefly describe the model distribution adopted for different binary orbital parameters based on the two models.
3.1.1 Mass of Primary, $m_{1}$
For the mass of primary, $m_{1}$, 2018AJ....156..257S assigned a fixed value of $m_{1}=0.8M_{\odot}$ because most stars in UMi are extremely old and are supposed to be located along the red-giant branch (RGB), where the masses are distributed in a very narrow range. However, we aim to adopt a more realistic mass distribution. To achieve this, we sample masses from an isochrone given by the PARSEC tracks [^3] http://stev.oapd.inaf.it/cgi-bin/cmd (2012MNRAS.427..127B). We select PARSEC version 1.2S (2014MNRAS.444.2525C; 2015MNRAS.452.1068C; 2018MNRAS.476..496F) rather than the latest 2.0 to obtain an isochrone with a lower metallicity of [M/H]=2.2=-2.2, which is the lower limit of the PARSEC tracks but corresponds to the median metallicity of our sample. The age is set to 10 Gyr, reflecting the average age of UMi (2002AJ....123.3199C). We clip the faint end of the isochrone at an absolute magnitude limit of $M_{G}>1.46$ mag, which is derived from the faintest star in our data with an apparent magnitude of $m_{G}=20.92$ and the distance to UMi. Effectively, the magnitude limit already excludes the fainter, compacter parts such as main-sequence stars and dwarfs, which suggests a surface gravity limit of $\log g<3.0$ Additionally, we also remove the horizontal branch, as we have done with the real data.
We randomly assign the initial mass of each mock star using the initial mass function derived by 2002Sci...295...82K. Given the initial mass, we then extract the corresponding current mass, surface gravity, and G-band magnitude from the isochrone. To ensure consistency with the observational data, we apply a magnitude selection function to enforce the same G-band magnitude distribution in the mock sample as observed in our real dataset. The mass is distributed around $0.8M_{\odot}$, with outliers contributing only a small fraction. In contrast, surface gravity spans a much wider range from 0.9 to over 3.0, which differs significantly from the assumption made by 2018AJ....156..257S of a constant value of $1.0$. This broader range is mainly due to the depth of DESI observation, which includes fainter/compacter objects. Since surface gravity plays an important role when generating the orbital period, it is necessary to adopt the more reliable values obtained from the isochrone.
3.1.2 Mass Ratio, q
@@
1991A&A...248..485D takes a normal distribution of $q$ as
@@
Here we also consider the excess possibility of twin binaries, which has been reported to have a clear narrow peak in the probability distribution of the mass ratio at $q>0.95$ for close binaries (2000A&A...360..997T; 2006ApJ...639L..67P; 2010ApJS..190....1R). The excess probability is defined as
@@
We present only the relevant parts of the distribution applicable to our study. The reader can refer to 2017ApJS..230...15M for detailed information on the complete formulations for other primary mass ranges.
3.1.3 Period, P
The adopted orbital period distribution most significantly affects the binary fraction constraints in our analysis (2013ApJ...779..116M; 2017AJ....153..254S, 2018AJ....156..257S). 1991A&A...248..485D model the orbital period distribution with a log-normal form
@@ 3.1.3 Period, P
We adopt $\mu_{\log P}=4.8$ and $\sigma_{\log P}=2.3$, where the period is in units of day. The adopted mean and standard deviation values provide the best fit to F7 to G9 type stars in our solar neighborhood, as found by 1991A&A...248..485D. We also investigated different values from several other studies, including 2010ApJS..190....1R, 1992ApJ...396..178F, and 2011MNRAS.417.1702M. A larger mean value of the period decreases the inferred binary fraction, whereas a larger standard deviation increases it. Specifically, a larger standard deviation leads to a higher fraction of short-period binaries, which have a significant effect on the LOSV variabilities (see more details in Section 3.3 below). In contrast, medium- to long-period binaries have minimal impact on the results due to the fact that our UMi member star observations are at most spanned by one year, which cannot provide strong constraints on medium- or long-period binaries. In this work, we focus on the values from 1991A&A...248..485D which have been demonstrated as the best model among the works considered above, as supported by 2018AJ....156..257S.
Using Kepler’s third law, we can constrain the minimum and maximum periods based on the semi-major axis, $a$. We adopt the same expression for the maximum semi-major axis as used by 2018AJ....156..257S, which is considered as the mean free path for stellar collision,
@@
where $\sigma_{v}=8.0$ km-1 is the velocity dispersion of UMi (2018AJ....156..257S), $t=10$ Gyr is the average age of its member stars (2002AJ....123.3199C), and $\lambda=0.2344$ pc-3 is the stellar number density, which is derived by the central luminosity density $L=0.006L_{\odot}$ pc-3 (1998ARA&A..36..435M) and the assumption that the average mass of the stars in UMi is $0.4M_{\odot}$, applying the mass-luminosity relation $L/L_{\odot}\propto(M/M_{\odot})^{4}$. The corresponding maximum limit of the period is around $10^{7}$ days.
The lower limit of the semi-major axis is considered as the radius of the Roche lobe of the primary, which depends on the mass ratio $q$. The analytical formula of the approximate Roche radius is given by 1983ApJ...268..368E,
@@
On the other hand, 2017ApJS..230...15M demonstrates a more complicated distribution for the orbital period as a function of the primary mass $m_{1}$. We do not repeatedly show the form here, and the readers can check 2017ApJS..230...15M for details. These values are all based on their measurements, combining numerous observations. The equations presented above are valid for systems with mass ratios of $q>0.3$. For systems with smaller mass ratios $0.1<q<0.3$, the overall distribution remains qualitatively similar, exhibiting only minor amplitude variations that do not impact our sampling. Therefore, to simplify the simulation, we adopt the same analytical functions to describe systems across the entire range of mass ratios $0.1<q<1.0$, as the variations are not significant enough to be distinct.
3.1.4 Eccentricity, e
2018AJ....156..257S choose a uniform distribution to simplify the model since the eccentricity only has a small effect on the observed binary fraction.
@@ 3.1.4 Eccentricity, e
2017ApJS..230...15M investigated a power law distribution for the eccentricity,
@@ 3.1.5 Inclination, i
3.1.6 Argument of Periastron, $\omega$
@@
3.1.7 True Anomaly, $\theta$
@@
3.2 Monte Carlo Simulations
@@
    Add Gaussian Noise: We add Gaussian random noise to the LOSV of both binary and non-binary stars (with the LOSV centered at zero for non-binary stars), using the quadrature sum of the measurement uncertainty, $\sigma_{\rm{obs}}$, and the fixed systematic error floor, $\sigma_{\rm{sys}}$, as the standard deviation, $\sigma_{v_{\rm{los}}}=\sqrt{\sigma_{{\rm{obs}}}^{2}+\sigma_{\rm{sys}}^{2}}$. When $\sigma_{\rm{sys}}$ is not yet fixed in the analysis, it is treated as a variable parameter and adjusted accordingly during the simulation process (see Section 3.5 for further details).
@@
3.3 LOSV Variability, $\beta$
Following 2018AJ....156..257S, we can define the LOSV variation for each pair of observations as
@@ 3.3 LOSV Variability, $\beta$
where $v_{\mathrm{los}}$ is the LOSV and $\sigma_{v_{\mathrm{los}}}$ is the corresponding uncertainty. The subscripts $i$ and $j$ denote different observations of the same star. Thus, for a star with multiple observations, we calculate an $\alpha$ value for each unique pair of observations. Consequently, a star with $n$ observations contributes n(n1)/2n(n-1)/2 pairs with different time separations to the $\alpha$ distribution. Notably, before constraining the systematic error floor in our analysis, the uncertainties in the denominator include only the observational errors. Once the systematic error floor is determined, we repeat the analysis by incorporating it into the total uncertainty in quadrature. Further details on this procedure can be found in Section 3.5.
@@
and use this quantity in the subsequent Bayesian analysis (see Section 3.4).
@@
3.4 Likelihood Function
As described in Section 3.2, we derive the $\beta$ distributions for both the binary and non-binary mock star samples, as shown in Figure 4. The probability distributions of $\beta$ are constructed from all $10^{5}$ mock realizations and normalized to unity. The black histogram represents the $\beta$ distribution obtained from the observational data. By combining the binary and non-binary samples with varying fractions, we can infer the binary fraction that best reproduces the observed distribution.
@@ 3.4 Likelihood Function
We apply a Markov Chain Monte Carlo (MCMC) approach to estimate the binary fraction and its associated uncertainty. For this purpose, we use the python package emcee [^4] https://github.com/dfm/emcee (emcee), an affine-invariant ensemble sampler designed for Bayesian parameter estimation. In this analysis, we construct the posterior probability distribution by combining the likelihood function derived from the observed $\beta$ distribution with a uniform prior to the binary fraction. The final posterior distribution provides both the best-fit binary fraction and its credible interval, which we report as the 68% highest posterior density interval (HPDI).
3.5 Determination of Systematic Error Floor of LOSV
@@
To overcome this limitation, we adopt a grid-based strategy. We vary the systematic error floor $\sigma_{\rm sys}$ from 0.8 to 2.0 km s-1 in increments of 0.1 km s-1. For each value, we recalculate the corresponding $\beta$ distribution by adding Gaussian noise to the simulated LOSVs, using the total uncertainty, $\sigma_{v_{\rm{los}}}=\sqrt{\sigma_{{\rm{obs}}}^{2}+\sigma_{\rm{sys}}^{2}}$ as described in Step 3 of Section 3.2. We then perform the MCMC estimation of the binary fraction based on each recalculated $\beta$ distribution.
To identify the most suitable value of the systematic error floor, we compare the maximum log-likelihood values obtained from the MCMC samplers for each $\sigma_{\rm sys}$. Figure 5 presents the relative maximum log-likelihoods for different systematic error floor values, with the values normalized by subtracting the absolute maximum log-likelihood to highlight the relative differences. As shown in the figure, the maximum log-likelihood is achieved when the systematic error floor is $\sigma_{\rm sys}=1.2\ {\rm km\ s}^{-1}$, indicating that this value best characterizes the systematic error floor under the observing conditions of our sample. This value is also consistent with the values estimated in 2025arXiv250514787K. Accordingly, we adopt $\sigma_{\rm{sys}}=1.2\ \rm{km\ s}^{-1}$ as the fixed systematic error floor in our final analysis and use it to construct the posterior distribution for the binary fraction. For all of the analyses in the following sections, the binary fraction is the only free model parameter.
@@
4 Results
4.1 Binary Fraction in UMi
We derive the binary fraction of UMi using both 1991A&A...248..485D and 2017ApJS..230...15M models. The posterior probability distribution (PPD) is shown in Figure 6. From the 1991A&A...248..485D model, we obtain a binary fraction of $0.61^{+0.16}_{-0.20}$, while the 2017ApJS..230...15M model yields a result of $0.69^{+0.19}_{-0.17}$. The quoted $1\sigma$ uncertainties are derived from the 68% HPDI of the cumulative PPDs. Our results from the 2017ApJS..230...15M model agree well with the previous work done by 2018AJ....156..257S, who found a best-fit binary fraction of $0.78^{+0.09}_{-0.08}$. Although their study also followed the 1991A&A...248..485D framework, our approach integrates a stellar isochrone in the modeling procedure to generate quantities including the mass of the primary star and surface gravity for each star (see Section 3.1 for details), providing a more physically motivated representation of the stellar population in UMi.
Figure 6: The top and bottom panels show the best constraint PPDs of binary fractions based on the 1991A&A...248..485D and 2017ApJS..230...15M models, respectively. The blue dashed line indicates the median and the red dashed lines show the $1\sigma$ range.
We further subdivide our sample into metal-rich and metal-poor subsamples based on their metallicities. For stars with multiple observations, we obtain their metallicity from the stacked spectra (see Section 2.2). We divide stars into metal-rich and metal-poor populations based on their metallicity using a threshold of $\mathrm{Fe}/\mathrm{H}]=-2.14$. This value, 2.14-2.14, corresponds to the median metallicity of the full sample; thereby our metal-rich ($\mathrm{Fe}/\mathrm{H}]>-2.14$) and metal-poor ($\mathrm{Fe}/\mathrm{H}]<-2.14$) subsamples contain equal number of stars. Explicitly, we have a metal-rich sample of 335 stars with a total of 1,024 observations at different epochs. The metal-poor sample also has 335 stars, with a total of 1,123 observations.

### v2

@@ III Methodology
Motivated by the method developed in Spencer et al. (2017a) and S18, we introduce several improvements to achieve a more realistic constraint on the binary fraction and a more robust estimation of its uncertainty. Specifically, we refine both the modeling of orbital parameters and the construction of the likelihood function. The detailed methodological differences will be described in the following subsections.
@@
III.1 Binary Orbital Parameters
@@
We adopt Equation 1 to simulate the LOSV at different epochs of a given star (see more details in Section III.2 below). To achieve this goal, we have to know the probability distribution of the seven orbital parameters, i.e., $m_{1}$, $q$, $P$, $e$, $i$, $\omega$, and $\theta$.
Among these parameters, $q$, $m_{1}$, $e$, and $P$ are intrinsic, while the other three angles of $i$, $\theta$, and $\omega$ depend on the observer’s perspective. We adopt the distributions of the intrinsic parameters ($q$, $e$, $P$) from two different models developed by Duquennoy and Mayor (1991)(hereafter DM91) and Moe and Di Stefano (2017)(hereafter MS17), which constrain the binary orbital parameter distributions based on solar neighborhood observations. The DM91 model is widely used in many previous studies. It provides independent distributions for each orbital parameter and is based on solar-like stars, making it straightforward to apply. It is important to note that our member stars in UMi are all giant stars with metallicity of $\mathrm{[Fe/H]}<-1$, significantly lower than those in the solar neighborhood. This means that by adopting the DM91 model, we are assuming that the binary properties of giants in dwarf galaxies are similar to those of solar-like stars in the solar neighborhood.
@@
We refer the readers to the original DM91 and MS17 papers for more details, and in the following, we briefly describe the model distribution adopted for different binary orbital parameters based on the two models.
III.1.1 Mass of Primary, $m_{1}$
For the mass of primary, $m_{1}$, S18 assigned a fixed value of $m_{1}=0.8M_{\odot}$ because most stars in UMi are extremely old and are supposed to be located along the red-giant branch (RGB), where the masses are distributed in a very narrow range. However, we aim to adopt a more realistic mass distribution. To achieve this, we sample masses from an isochrone given by the PARSEC tracks [^3] http://stev.oapd.inaf.it/cgi-bin/cmd (Bressan et al., 2012). We select PARSEC version 1.2S (Chen et al., 2014, 2015; Fu et al., 2018) rather than the latest 2.0 to obtain an isochrone with a lower metallicity of [M/H]=2.2=-2.2, which is the lower limit of the PARSEC tracks but corresponds to the median metallicity of our sample. The age is set to 10 Gyr, reflecting the average age of UMi (Carrera et al., 2002). We clip the faint end of the isochrone at an absolute magnitude limit of $M_{G}>1.46$ mag, which is derived from the faintest star in our data with an apparent magnitude of $m_{G}=20.92$ and the distance to UMi. Effectively, the magnitude limit already excludes the fainter, compacter parts such as main-sequence stars and dwarfs, which suggests a surface gravity limit of $\log g<3.0$ Additionally, we also remove the horizontal branch, as we have done with the real data.
We randomly assign the initial mass of each mock star using the initial mass function derived by Kroupa (2002). Given the initial mass, we then extract the corresponding current mass, surface gravity, and G-band magnitude from the isochrone. To ensure consistency with the observational data, we apply a magnitude selection function to enforce the same G-band magnitude distribution in the mock sample as observed in our real dataset. The mass is distributed around $0.8M_{\odot}$, with outliers contributing only a small fraction. In contrast, surface gravity spans a much wider range from 0.9 to over 3.0, which differs significantly from the assumption made by S18 of a constant value of $1.0$. This broader range is mainly due to the depth of DESI observation, which includes fainter/compacter objects. Since surface gravity plays an important role when generating the orbital period, it is necessary to adopt the more reliable values obtained from the isochrone.
III.1.2 Mass Ratio, q
@@
DM91 takes a normal distribution of $q$ as
@@
Here we also consider the excess possibility of twin binaries, which has been reported to have a clear narrow peak in the probability distribution of the mass ratio at $q>0.95$ for close binaries (Tokovinin, 2000; Pinsonneault and Stanek, 2006; Raghavan et al., 2010). The excess probability is defined as
@@
We present only the relevant parts of the distribution applicable to our study. The reader can refer to MS17 for detailed information on the complete formulations for other primary mass ranges.
III.1.3 Period, P
The adopted orbital period distribution most significantly affects the binary fraction constraints in our analysis (Minor, 2013; Spencer et al., 2017a, S18). DM91 model the orbital period distribution with a log-normal form
@@ III.1.3 Period, P
We adopt $\mu_{\log P}=4.8$ and $\sigma_{\log P}=2.3$, where the period is in units of day. The adopted mean and standard deviation values provide the best fit to F7 to G9 type stars in our solar neighborhood, as found by DM91. We also investigated different values from several other studies, including Raghavan et al. (2010), Fischer and Marcy (1992), and Marks and Kroupa (2011). A larger mean value of the period decreases the inferred binary fraction, whereas a larger standard deviation increases it. Specifically, a larger standard deviation leads to a higher fraction of short-period binaries, which have a significant effect on the LOSV variabilities (see more details in Section III.3 below). In contrast, medium- to long-period binaries have minimal impact on the results due to the fact that our UMi member star observations are at most spanned by one year, which cannot provide strong constraints on medium- or long-period binaries. In this work, we focus on the values from DM91 which have been demonstrated as the best model among the works considered above, as supported by S18.
Using Kepler’s third law, we can constrain the minimum and maximum periods based on the semi-major axis, $a$. We adopt the same expression for the maximum semi-major axis as used by S18, which is considered as the mean free path for stellar collision,
@@
where $\sigma_{v}=8.0$ km-1 is the velocity dispersion of UMi (S18), $t=10$ Gyr is the average age of its member stars (Carrera et al., 2002), and $\lambda=0.2344$ pc-3 is the stellar number density, which is derived by the central luminosity density $L=0.006L_{\odot}$ pc-3 (Mateo, 1998) and the assumption that the average mass of the stars in UMi is $0.4M_{\odot}$, applying the mass-luminosity relation $L/L_{\odot}\propto(M/M_{\odot})^{4}$. The corresponding maximum limit of the period is around $10^{7}$ days.
The lower limit of the semi-major axis is considered as the radius of the Roche lobe of the primary, which depends on the mass ratio $q$. The analytical formula of the approximate Roche radius is given by Eggleton (1983),
@@
On the other hand, MS17 demonstrates a more complicated distribution for the orbital period as a function of the primary mass $m_{1}$. We do not repeatedly show the form here, and the readers can check Moe and Di Stefano (2017) for details. These values are all based on their measurements, combining numerous observations. The equations presented above are valid for systems with mass ratios of $q>0.3$. For systems with smaller mass ratios $0.1<q<0.3$, the overall distribution remains qualitatively similar, exhibiting only minor amplitude variations that do not impact our sampling. Therefore, to simplify the simulation, we adopt the same analytical functions to describe systems across the entire range of mass ratios $0.1<q<1.0$, as the variations are not significant enough to be distinct.
III.1.4 Eccentricity, e
S18 choose a uniform distribution to simplify the model since the eccentricity only has a small effect on the observed binary fraction.
@@ III.1.4 Eccentricity, e
MS17 investigated a power law distribution for the eccentricity,
@@ III.1.5 Inclination, i
III.1.6 Argument of Periastron, $\omega$
@@
III.1.7 True Anomaly, $\theta$
@@
III.2 Monte Carlo Simulations
@@
    Add Gaussian Noise: We add Gaussian random noise to the LOSV of both binary and non-binary stars (with the LOSV centered at zero for non-binary stars), using the quadrature sum of the measurement uncertainty, $\sigma_{\rm{obs}}$, and the fixed systematic error floor, $\sigma_{\rm{sys}}$, as the standard deviation, $\sigma_{v_{\rm{los}}}=\sqrt{\sigma_{{\rm{obs}}}^{2}+\sigma_{\rm{sys}}^{2}}$. When $\sigma_{\rm{sys}}$ is not yet fixed in the analysis, it is treated as a variable parameter and adjusted accordingly during the simulation process (see Section III.5 for further details).
@@
III.3 LOSV Variability, $\beta$
Following Spencer et al. (2018), we can define the LOSV variation for each pair of observations as
@@ III.3 LOSV Variability, $\beta$
where $v_{\mathrm{los}}$ is the LOSV and $\sigma_{v_{\mathrm{los}}}$ is the corresponding uncertainty. The subscripts $i$ and $j$ denote different observations of the same star. Thus, for a star with multiple observations, we calculate an $\alpha$ value for each unique pair of observations. Consequently, a star with $n$ observations contributes n(n1)/2n(n-1)/2 pairs with different time separations to the $\alpha$ distribution. Notably, before constraining the systematic error floor in our analysis, the uncertainties in the denominator include only the observational errors. Once the systematic error floor is determined, we repeat the analysis by incorporating it into the total uncertainty in quadrature. Further details on this procedure can be found in Section III.5.
@@
and use this quantity in the subsequent Bayesian analysis (see Section III.4).
@@
III.4 Likelihood Function
As described in Section III.2, we derive the $\beta$ distributions for both the binary and non-binary mock star samples, as shown in Figure 4. The probability distributions of $\beta$ are constructed from all $10^{5}$ mock realizations and normalized to unity. The black histogram represents the $\beta$ distribution obtained from the observational data. By combining the binary and non-binary samples with varying fractions, we can infer the binary fraction that best reproduces the observed distribution.
@@ III.4 Likelihood Function
We apply a Markov Chain Monte Carlo (MCMC) approach to estimate the binary fraction and its associated uncertainty. For this purpose, we use the python package emcee [^4] https://github.com/dfm/emcee (Foreman-Mackey et al., 2013), an affine-invariant ensemble sampler designed for Bayesian parameter estimation. In this analysis, we construct the posterior probability distribution by combining the likelihood function derived from the observed $\beta$ distribution with a uniform prior to the binary fraction. The final posterior distribution provides both the best-fit binary fraction and its credible interval, which we report as the 68% highest posterior density interval (HPDI).
III.5 Determination of Systematic Error Floor of LOSV
@@
To overcome this limitation, we adopt a grid-based strategy. We vary the systematic error floor $\sigma_{\rm sys}$ from 0.8 to 2.0 km s-1 in increments of 0.1 km s-1. For each value, we recalculate the corresponding $\beta$ distribution by adding Gaussian noise to the simulated LOSVs, using the total uncertainty, $\sigma_{v_{\rm{los}}}=\sqrt{\sigma_{{\rm{obs}}}^{2}+\sigma_{\rm{sys}}^{2}}$ as described in Step 3 of Section III.2. We then perform the MCMC estimation of the binary fraction based on each recalculated $\beta$ distribution.
To identify the most suitable value of the systematic error floor, we compare the maximum log-likelihood values obtained from the MCMC samplers for each $\sigma_{\rm sys}$. Figure 5 presents the relative maximum log-likelihoods for different systematic error floor values, with the values normalized by subtracting the absolute maximum log-likelihood to highlight the relative differences. As shown in the figure, the maximum log-likelihood is achieved when the systematic error floor is $\sigma_{\rm sys}=1.2\ {\rm km\ s}^{-1}$, indicating that this value best characterizes the systematic error floor under the observing conditions of our sample. This value is also consistent with the values estimated in Koposov et al. (2025). Accordingly, we adopt $\sigma_{\rm{sys}}=1.2\ \rm{km\ s}^{-1}$ as the fixed systematic error floor in our final analysis and use it to construct the posterior distribution for the binary fraction. For all of the analyses in the following sections, the binary fraction is the only free model parameter.
@@
IV Results
IV.1 Binary Fraction in UMi
We derive the binary fraction of UMi using both DM91 and MS17 models. The posterior probability distribution (PPD) is shown in Figure 6. From the DM91 model, we obtain a binary fraction of $0.61^{+0.16}_{-0.20}$, while the MS17 model yields a result of $0.69^{+0.19}_{-0.17}$. The quoted $1\sigma$ uncertainties are derived from the 68% HPDI of the cumulative PPDs. While Olszewski et al. (1996) reported an early, highly uncertain estimate of the binary frequency near one-year periods in UMi, the expanded multi-epoch LOSV dataset and improved binary population modeling in this work yield substantially more robust constraints on the total binary fraction. Our results from the MS17 model agree well with the previous work done by S18, who found a best-fit binary fraction of $0.78^{+0.09}_{-0.08}$. Although their study also followed the DM91 framework, our approach integrates a stellar isochrone in the modeling procedure to generate quantities including the mass of the primary star and surface gravity for each star (see Section III.1 for details), providing a more physically motivated representation of the stellar population in UMi. Our results exhibit larger uncertainties than those reported by S18, primarily due to differences in the definition of the LOSV variability, $\beta$. As discussed in III.3, the approach adopted by S18 effectively amplifies the contribution of stars with a larger number of observations. Consequently, multiple $\beta$ values from the same star enter the likelihood function, which increases the apparent statistical weight of these stars and leads to an underestimation of the overall uncertainties. On the other hand, our Equation 20 forces the same weight from each individual star.
Figure 6: The top and bottom panels show the best constraint PPDs of binary fractions based on the DM91 and MS17 models, respectively. The blue dashed line indicates the median and the red dashed lines show the $1\sigma$ range.
We further subdivide our sample into metal-rich and metal-poor subsamples based on their metallicities. For stars with multiple observations, we obtain their metallicity from the stacked spectra (see Section II.2). We divide stars into metal-rich and metal-poor populations based on their metallicity using a threshold of $\mathrm{Fe}/\mathrm{H}]=-2.14$. This value, 2.14-2.14, corresponds to the median metallicity of the full sample; thereby our metal-rich ($\mathrm{Fe}/\mathrm{H}]>-2.14$) and metal-poor ($\mathrm{Fe}/\mathrm{H}]<-2.14$) subsamples contain equal number of stars. Explicitly, we have a metal-rich sample of 335 stars with a total of 1,024 observations at different epochs. The metal-poor sample also has 335 stars, with a total of 1,123 observations.